OpenStudy (anonymous):

What is the midpoint of (–2, 0) and (4, –2)? A.(0, 0) B.(2, –2) C.(1, –1) D.(–6, 2) I Have Always Been Bad At Math Can Anyone Answer This Question For Me ?

OpenStudy (jlastino):

The mid point is the point that is exactly at the middle of the two given points Add the two values of x and divide the sum by 2 -2+4 = 2 2/2 = 1 So the x coordinate would be 1 Add the two values of y and divide the sum by 2 0+-2= 0-2 = -2 -2/2 = -1 So the midpoint will be (1,-1)